OpenStudy (anonymous):

multiply. -[(-3)(x-y)]

OpenStudy (anonymous):

[(-3)(x-y)] 3(x+y) -[3x+3y] =3x+3y

OpenStudy (anonymous):

does this help?

OpenStudy (anonymous):

Examples: |4| = 4 |-4| = 4 |4+3| = 7 |-4-3| = 7 |3-4| = 1 -|4| = -4 -|-4| = -4 here's the idea